OpenStudy (anonymous):

7 3/10 + 6 1/3 + 2 7/10 A. 15 1/3 B. 16 1/3 C. 16 11/23 D. 16 5/6

OpenStudy (alexandervonhumboldt2):

first we need to make denominators equal: it is good to make denomitaor 30 in your case: 7 9/30 + 6 10/30 + 2 21/30=15 40/30=16 10/30=16 1/3

OpenStudy (anonymous):

Sorry, I was brb, thanks! :)